OpenStudy (anonymous):

what would x^2-16/x^2-4x be in simplest form or can someone explain how to do it

OpenStudy (blurbendy):

can you draw the equation please

OpenStudy (anonymous):

\[ \frac{ x^2-16}{ x^2-4x }\]

OpenStudy (blurbendy):

okay let's break it down. how can you rewrite the numerator (top part) ?

OpenStudy (anonymous):

(x+8)(x-8)

OpenStudy (blurbendy):

close, it's actually (x-4)(x+4)

OpenStudy (blurbendy):

now, how we can rewrite the denominator (bottom part)

OpenStudy (anonymous):

(x-2)(x+2) but what about the extra x

OpenStudy (blurbendy):

not quite. simplify it like this: x(x-4) since x^2 - 4x both share an x, we're just going to take it out.

OpenStudy (blurbendy):

now, what does the whole expression look like numerator and denominator

OpenStudy (anonymous):

(x+4)/x right?

OpenStudy (blurbendy):

good, so you cancelled out the (x-4)?

OpenStudy (anonymous):

yes now its done right

OpenStudy (blurbendy):

(x+4) / x can be written as x/x + 4/x = 1 + 4/x that's as far as you can simplify it

OpenStudy (anonymous):

good i hate math and this packet is killing me :)

OpenStudy (blurbendy):

lol you did good though. keep it up!

OpenStudy (anonymous):

hahah thank you